A journalist would like to report on the traffic–related tardiness incurred by employees

in a certain city per year. To do this, he asked 25 employees and found out that they were

late for an average of 8.5 times a year due to traffic. The sample standard deviation was found to be in 1.5 times.


1. At a confidence level of 90 percent, determine the confidence interval that will

estimate the population mean.


2. At a confidence level of 95 percent, determine the confidence interval that will

estimate the average number of traffic-related tardiness per year by employees in

a city.


3. Which between the answers to items a and b gave a wider margin of error? What

does this say about the relationship between the confidence level and the margin

of error?


1
Expert's answer
2021-06-22T13:48:46-0400

1.

The critical value for "\alpha=0.10," and "df=n-1=25-1=24" degrees of freedom is "t_c=1.710882."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(8.5-1.710882\times\dfrac{1.5}{\sqrt{25}}, 8.5+1.710882\times\dfrac{1.5}{\sqrt{25}})"

"=(7.987, 9.013)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "7.987<\mu<9.013," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(7.987, 9.013)."


2.

The critical value for "\alpha=0.05," and "df=n-1=25-1=24" degrees of freedom is "t_c=2.063899."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(8.5-2.063899\times\dfrac{1.5}{\sqrt{25}}, 8.5+2.063899\times\dfrac{1.5}{\sqrt{25}})"

"=(7.881, 9.119)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "7.881<\mu<9.119," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(7.881, 9.119)."


3.

The 90% confidence interval for the population mean is "7.987<\mu<9.013."

The 95% confidence interval for the population mean is "7.881<\mu<9.119."

The higher the confidence level, the wider the confidence interval is (if everything else is equal).

As the confidence level increases, the margin of error increases.